Find the general/particular solution of the following Differential Equations

(Exact D.E)


1.) ds/dt=(s²Sin t)/(2 Cos t)




1
Expert's answer
2021-09-08T18:36:24-0400

"\frac{ds}{dt}=\frac{s^2sint}{2cost}\\
2\frac{ds}{s^2}=tant\space dt\\
\text{Integrating both side,}\\
-2s^{-1}=ln|sect|+c\\
\text{This is required general solution.}"